OpenStudy (anonymous):

the difference of two numbers is 2. if the sum of their reciprocals is 9/40, find the two numbers

OpenStudy (misty1212):

\[x-y=2\\ \frac{1}{x}+\frac{1}{y}=\frac{9}{40}\] is how i would begin it

OpenStudy (misty1212):

or maybe \[\frac{1}{x}+\frac{1}{x+2}=\frac{9}{40}\]

OpenStudy (misty1212):

need more help with this?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

do you cross multiply?

OpenStudy (misty1212):

OK

OpenStudy (misty1212):

we can add them

OpenStudy (triciaal):

before you cross multiply you need to simplify the left hand side

OpenStudy (misty1212):

\[\frac{x+2+x}{x(x+2)}=\frac{9}{40}\]

OpenStudy (misty1212):

or just \[\frac{2x+2}{x(x+2)}=\frac{9}{40}\]

OpenStudy (solomonzelman):

x-y=2 x=y+2 and x-2=y, NOT x+2=y

OpenStudy (solomonzelman):

I think you might have made a mistake

OpenStudy (misty1212):

probably

OpenStudy (solomonzelman):

yes, check your work, please

OpenStudy (misty1212):

\[\frac{1}{x}+\frac{1}{x-2}=\frac{9}{40}\] might work better

OpenStudy (solomonzelman):

good. Won't interrupt any longer:) Do your work:D

OpenStudy (anonymous):

i got 80x+80 ------- 9x^2-18x

OpenStudy (misty1212):

of course is actually makes no difference at all

OpenStudy (misty1212):

lets add \[\frac{x-2+x}{x(x-2)}=\frac{9}{40}\] or \[\frac{2x-2}{x(x-2)}=\frac{9}{40}\]

OpenStudy (misty1212):

then go ahead and "cross multiply" like you said, get \[80x-80=9x^2-18x\]

OpenStudy (misty1212):

now you have a quadratic equation to solve

OpenStudy (misty1212):

\[9x^2-18x-80x+80=0\\ 9x^2-98x+80=0\] which you can factor

OpenStudy (misty1212):

\[(9x-8)(x-10)=0\] et

OpenStudy (misty1212):

you get two different numbers for \(x\) which leads to two different y values

OpenStudy (misty1212):

i would like to point out that if you solve \[\frac{1}{x}+\frac{1}{x+2}=\frac{4}{90}\] you get the same solutions no one can tell \(x\) from \(y\) in this question

OpenStudy (anonymous):

i got x=10 & y=8